OpenStudy (anonymous):

5+3+2= 151022 9+2+4= 183612 8+6+3= 482466 5+4+5= 202541 then 7+2+5 = ........?

OpenStudy (anonymous):

1435

OpenStudy (anonymous):

last two digits?

OpenStudy (anonymous):

wait.....

OpenStudy (anonymous):

I solved upto that ^^

OpenStudy (anonymous):

47

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

sum of the above two minus the second number in the question (15+10)-3=22 (18+36)-2=52 (48+24)-6=66 (20+25)-4=41

OpenStudy (anonymous):

in the second one, that's 12^^

OpenStudy (anonymous):

5*3=15, 5*2=10, 5*(3+2)=25-3=22 (151022) 9*2=18, 9*4=36, 9*(2+4)=54-2=52 (183652) 8*6=48, 8*3=24, 8*(6+3)=72-6=66 (482466) 5*4=20, 5*5=25, 5*(4+5)=45-4=41 (202541) hence ; 7*2=14, 7*5=35, 7*(2+5)=49-2=47 (143547)